Using the thesis-me Class File

R W Kaszeta

8 May 1997

This document describes the features and usage of the thesis-me class file for typesetting theses according to the guidelines set by the Graduate School in their document, “Preparation of the Doctoral Thesis/Project”.

What this class file does:

This class file provides a modified “book” class for L^{A}TEX 2ε, so most anything that would work for a “book” will work in your thesis. The changes from the standard “book” class are:

1. The default paragraph indent is set to 0 points.

2. The default paragraph separation is set to 10 points.

3. The page margins are set at 1.5 inches on the left, 1 inches elsewhere.

4. The linespacing is set to 1.5.

5. The section, subsection, and subsubsection headers have been changed so the aren’t as large as the L^{A}TEX 2εdefaults.

6. The \maketitle macro has been changed to produce a signature page, a title page, and a copyright page of the correct format.

7. Page numbering is set so that the title pages are unnumbered, front matter is numbered with roman numerals, and all subsequent pages are number with arabic numerals.

8. Two-sided typesetting is disabled.

9. abstract, dedication, and acknowledgments environments are defined for use in the front matter.

10. The endnotes document class option redefines the \footnote macros to place footnotes at the end of the chapter instead of the bottom of the page.

What this class doesn’t do:

This package only makes the changes required by the Graduate School, it doesn’t make changes for user preferences. Since the package is basically compatible with the standard L^{A}TEX 2ε classes, the standard L^{A}TEX 2ε customization methods and packages can be used.

Using this class:

The first step to using this class file is to tell L^{A}TEX 2εthat you want to use the thesis-me class:

\documentclass[]{thesis-me}

which loads the class file.

Then, you must provide the following information:

\Author{Author Name}
\Title{Thesis Title}
\Month{Month of Thesis Publication}
\Year{Year of Thesis Publication}
\Adviser{Adviser Name}

and if you are a Masters student you should also insert

\Degree{MASTER OF SCIENCE}
\degreelevel{masters}

To produce the standard signature, title, and copyright pages, simply start your document

with a \maketitle.

After the title pages, you may wish to include an abstract and dedication, these are done with the abstract and dedication environments.

All “front matter” (everything after the title page and before the first chapter) should be preceded by the macro \frontmatter which resets the page counter and tells L^{A}TEX 2ε to number the pages with roman numerals.

Useful things to include in the front matter include

\tableofcontents
\listoffigures
\listoftables

After the front matter you are ready to begin the main document, so you enter a \mainmatter before the first chapter to tell L^{A}TEX 2ε to again reset the page counter and start numbering pages with arabic numerals.

After all the chapters, you enter the macro \appendix before the appendix chapters so that \chapter will start labelling the chapters as appendices.

The \backmatter macro can be used after the appendices for further chapters that you don’t want titled as “chapters”, such as an index.

A sample .tex file for this class is provided in /stage/share/tex.local/thesis-me-template.
tex.

Line Spacing:

This class defaults to 1.5 spacing for the document. If you would like to doublespace the thesis instead, simply insert a \doublespacing at the beginning of your document.

If you have environments like quotations, or entire listings like your bibliography which you would like singlespaced, you can include them inside a singlespace environment.

Endnotes

If you would like to use endnotes instead of footnotes, the first step is to tell the document class you want endnotes:

\documentclass[endnotes]{thesis-me}

Then, instead of using \footnote{}, you use \endnote instead. At the end of each chapter (or wherever you want the endnotes to appear) you can list all of the unprinted endnotes

using \showendnotes.

User Configuration:

In order to allow simple user configuration of this package without needing to modify thesis-me.cls, the class file loads in the tex file thesis-me.cfg (stuck someplace where your TEXINPUTS setting will find it) after it has done all of it’s configuration. This is a useful place to override the thesis-me default settings and provide macros you always wish to have handy when typesetting theses.

Other Useful Macro Packages

There are a number of additional macro packages that are useful for typesetting your thesis. All of these are documented on the MEnet TEX documentation page.

Package: Function:
graphicx Inclusion of graphics

rotating Allows landscape tables and figures, and rotating of graphical

material.

nomencl Provides an easy interface for creating nomenclature lists.
algorithms Provides an environment for typing programs and environments

Questions? Problems? Suggestions?

Contact kaszeta@me.umn.edu.

Appendix: A sample .tex file

\documentclass[12pt]{thesis-me}
\Author{Author Name}
\Title{Thesis Title}
\Month{Month of Thesis Publication}
\Year{Year of Thesis Publication}
\Adviser{Adviser Name}
\begin{document}
\maketitle
\frontmatter % stuff between the title and start of the document
\begin{dedication}
\end{dedication}
\newpage
\begin{abstract}
\end{abstract}
\newpage
\tableofcontents
\mainmatter % the main part of the document
\chapter{}
\chapter{}
\appendix % start labelling chapters as appendices
\chapter{}
\chapter{}
\end{document}